# SeekRBP: Leveraging Sequence-Structure Integration with Reinforcement Learning for Receptor-Binding Protein Identification

### 05 Mar 2026 v1

@@ 1 Introduction
Figure 1: Overview of the Multi-Armed Bandit (MAB)-driven dynamic negative sampling pipeline. A UCB1-based multi-armed bandit strategy iteratively selects informative hard negatives from a large pool and updates their utility using EL2N rewards after each training round. Sequence and structure representations are adaptively fused and fed into a classifier to predict RBP versus non-RBP.
@@ 2.2.1 Dual-branch protein encoding
The sequence branch uses ESM2 (Lin et al., 2023), a pre-trained protein language model, to extract representations from input protein sequences. Given an input sequence, ESM2 produces a high-dimensional embedding for each residue position. The embeddings are then processed by a Transformer encoder to obtain 1D sequence hidden representations $\mathbf{Z}_{1d}$ (see Supplementary 1.1).
@@
The structure branch encodes structure-derived features obtained by first predicting three-dimensional structures with ColabFold (Mirdita et al., 2022) and then extracting structure-aware representations using Saprot (Su et al., 2023). Specifically, ColabFold generates a predicted 3D structure for each input sequence, which is then processed by Saprot to produce residue-level structural features. These features are fed into a Transformer encoder to obtain 3D structure representations $\mathbf{Z}_{3d}$ (see Supplementary 1.2).
@@ 2.3 Training strategy and evaluation metrics
During the training process, the selected negative samples and all positive samples constitute a training batch used to update the model parameters. Following this update, the model’s predictions are refreshed, and the reward signal, specifically the prediction discrepancy $S^{l}$ (Eqn. 2), is recomputed to reflect the current state of the model. The model is trained using the Adam optimizer with a learning rate of $10^{-3}$ for a maximum of $30$ epochs. This iterative loop continues until the performance gain plateaus or a maximum number of epochs is reached. At the end of each cycle, the utility estimates $Q_{i}(l)$ and selection counts $n_{i}(l)$ are updated, informing the sampling policy for the subsequent round. More details are presented in the Supplementary 1.3.
@@
To evaluate the performance of the model in the RBPs binary classification task, we adopt Precision, Recall, and F1-score as the primary evaluation metrics (see Supplementary1.4). Given that RBP is the primary objective of the test, all metrics are calculated and reported with RBP defined as positive. We also plot the receiver operating characteristic (ROC) curve to illustrate the diagnostic ability of the classifiers by mapping its true positive rate (model’s ability to correctly identify RBPs) against the false positive rate (frequency of errors in which non-RBPs are misclassified as RBPs).
@@ 3.1.1 Effect of sampling strategies
First, we investigated the impact of different sampling strategies on model performance. Fig. 3 A indicates that our bandit-based sampling strategy achieves the highest AUC, particularly in regions with low false positive rates. The w/o Exploration variant represents a degenerate form of the bandit-based strategy where the exploration term in Eq. 1 is removed. While this exploitation-only strategy improves upon uniform sampling, it is consistently outperformed by the full UCB-based approach, highlighting the importance of explicitly encouraging the exploration of under-sampled negatives. Furthermore, we observe that the disparity between sampling strategies is sensitive to the input modality (Fig. S1 in supplementary). Notably, when the model relies only on 1-D sequence features, bandit-based sampling yields an AUC improvement of over 5%, suggesting that robust negative sampling is particularly critical when structural information is absent.
@@ 3.1.2 Contribution of sequence and structure features
Next, we evaluated the relative contributions of sequence and structure features. As illustrated in Fig. 3 B, the model utilizing only sequence features achieves an AUC of 0.921, whereas the structure-only model yields a slightly lower AUC of 0.913. This suggests that while tructural features are commonly regarded as informative, sequence data may offer greater discriminative power in our identification task. Notably, the integration of both modalities results in the highest performance (AUC 0.9418), consistently outperforming single-modality approaches. One plausible reason is that the model can utilize the complementary nature of the two features: sequence data provides broad discriminative signals, while structural features help resolve ambiguous cases that are difficult to distinguish via sequence alone. Thus, multimodal integration is essential for maximizing the robustness and accuracy of RBP classification.
@@ 3.1.3 Effectiveness of feature fusion strategies
Finally, we evaluated the effectiveness of different strategies for fusing sequence and structure representations. As shown in Fig. 3 C, simple feature concatenation (Concat) yields the lowest performance (AUC = 0.914), indicating that simply stacking features offers limited gains. While additive fusion (Sum) provides a slight improvement (AUC = 0.921), these naive aggregation methods remain insufficient for fully capturing cross-modal complementarity. In contrast, more advanced strategies consistently enhance performance: gated additive fusion increases the AUC to 0.931 by adaptively reweighting features, while low-rank multiplicative fusion achieves an AUC of 0.926, highlighting the benefits of modeling nonlinear cross-modal interactions. Ultimately, our proposed AEFM method achieves the highest AUC of 0.941. By dynamically combining additive and multiplicative interactions, AEFM integrates sequence and structural information more effectively than any single fusion strategy.
Figure 3: ROC curve comparison of ablation study. A) sampling strategies. The w/o Exploration variant represents a degenerate form of the bandit-based strategy. B) different feature modalities. Sequence only uses sequence-derived features, structure only uses structure-derived features, and sequence combined with structure integrates both modalities. C) feature fusion strategies. Sum: additive fusion; Concat: concatenation; AEFM: adaptive expert fusion module.
@@ 3.2 Comparison with benchmark methods
We benchmarked SeekRBP against state-of-the-art approaches, including learning-based tools (PhANNs, PhageRBPdetection) and homology-based methods (BLASTp, Pharokka). As illustrated in Fig. 4, homology-based methods achieve high precision (>0.95>0.95); however, most existing tools are constrained by low recall, highlighting the challenge of identifying divergent RBP families. In contrast, SeekRBP achieves the highest recall of 0.629 while maintaining competitive precision, resulting in the highest F1-score (0.742). This performance suggests that our bandit-based sampling strategy effectively identifies informative negative examples, enabling the model to learn discriminative features without sacrificing generalization. Furthermore, to accommodate diverse application scenarios, we offer an alternative configuration of SeekRBP that prioritizes high precision (~0.95), as detailed in Supplementary Table S2. Thus, SeekRBP provides a robust balance between precision and recall, making it well-suited for the real-world applications.
Figure 4: Comparison with benchmark methods on phage RBP identification. Precision, Recall, and F1-score are reported for PhANNs, PhageRBPdetection, Blastp, Pharokka, and our method, showing overall improved performance of the proposed approach.
@@ 3.3 Case study: application on vibrio phages
We evaluated the biological relevance and generalizability of the proposed framework using an independent experimental dataset of Vibrio phages. This dataset was generated through wet-lab experimental work by our collaborators, Zhang et al., involving purification, culturing, and induced lytic procedures (Zhang et al., 2025), where RBPs were identified via manual inspection of 3D structural domains. Although an absolute ground truth remains elusive, we utilized these manually curated RBPs as the reference standard. In this analysis, we first identified RBPs using SeekRBP and subsequently compared host prediction performance using the predicted RBPs versus the manually curated set.
@@
As shown in Supplementary Fig. S2, the model predicts a larger number of RBPs than the observed set while recovering a substantial fraction of manually selected RBPs across different phage genomes. In addition to count-based comparisons, we performed a structure-based analysis using TM-scores to assess the structural plausibility of predicted RBPs not identified manually. We computed the average TM-score and variance for manually selected RBPs that were not identified by our framework (the Observed set) and RBPs identified by SeekRBP that were not present in manual annotations (the Predicted set). A TM-score greater than 0.5 indicates significant structural similarity. As shown in Fig. 5, the predicted RBPs exhibit a wider range of TM-scores than the manually selected ones; notably, most predicted RBPs have a TM-score exceeding 0.5. This suggests that the SeekRBP detects divergent RBP candidates with valid structural folds or functional interactions that may have been overlooked under specific experimental conditions.
Figure 5: TM-score comparison between predicted RBPs (Predicted) and manually selected RBPs (Observed) in the Vibrio dataset. The predicted RBPs show larger TM-scores range, possibly reflecting discovered divergent RBP families.
@@
Furthermore, we investigated whether the additionally predicted RBPs could enhance downstream host recognition. To this end, we compared the performance of the Observed RBP set against the Predicted set using two complementary validation approaches. First, we evaluated their utility in a machine learning-based phage-host interaction prediction task. Using ESM-2 embeddings and a standard MLP classifier within a 10-fold stratified cross-validation framework (Supplementary S1.6), the Predicted set achieved a higher mean AUC (0.737) than the Observed set (0.713). The standard deviation also decreased from 0.099 to 0.073, indicating improved model stability (Supplementary Fig. S3). Second, to ensure these predictive gains reflect true biological interactions, we employed AlphaFold3 to estimate the docking quality (interface pLDDT) between the RBPs and host outer membrane proteins (Supplementary S1.7). Consistent with the classification results, the Predicted set exhibited higher interface pLDDT scores, pointing to more stable binding interfaces. Together, these findings demonstrate that the additional RBPs identified by SeekRBP provide robust biological signals rather than introducing noise, effectively bridging gaps in sparse experimental annotations.

### 16 Mar 2026 v2

@@ 1 Introduction
Figure 1: Overview of the SeekRBP’s dynamic negative sampling pipeline. A UCB1-based multi-armed bandit strategy iteratively selects informative hard negatives from a large pool and updates their utility using EL2N rewards after each training round. Sequence and structure representations are adaptively fused and fed into a classifier to predict RBP versus non-RBP.
@@ 2.2.1 Dual-branch protein encoding
The sequence branch uses ESM2 (Lin et al., 2023), a pre-trained protein language model, to extract representations from input protein sequences. Given an input sequence, ESM2 produces a high-dimensional embedding for each residue position. The embeddings are then processed by a Transformer encoder to obtain 1D sequence hidden representations $\mathbf{Z}_{1d}$ (see Supplementary S1.1).
@@
The structure branch encodes structure-derived features obtained by first predicting three-dimensional structures with ColabFold (Mirdita et al., 2022) and then extracting structure-aware representations using Saprot (Su et al., 2023). Specifically, ColabFold generates a predicted 3D structure for each input sequence, which is then processed by Saprot to produce residue-level structural features. These features are fed into a Transformer encoder to obtain 3D structure representations $\mathbf{Z}_{3d}$ (see Supplementary S1.2).
@@ 2.2.2 Adaptive expert fusion module (AEFM)
Figure 3: ROC curve comparison of ablation study. A) sampling strategies. The w/o Exploration variant represents a degenerate form of the bandit-based strategy. B) different feature modalities: sequence only uses sequence-derived features; structure only uses structure-derived features; and sequence combined both modalities (denoted as Ours). C) feature fusion strategies: Sum: additive fusion; Concat: concatenation; Gated Additive: only use additive interaction path; Low-Rank Multiplied: only use multiplicative interaction path; Ours: adaptive expert fusion module.
@@ 2.3 Training strategy and evaluation metrics
During the training process, the selected negative samples and all positive samples constitute a training batch used to update the model parameters. Following this update, the model’s predictions are refreshed, and the reward signal, specifically the prediction discrepancy $S^{l}$ (Eqn. 2), is recomputed to reflect the current state of the model. The model is trained using the Adam optimizer with a learning rate of $10^{-3}$ for a maximum of $30$ epochs. This iterative loop continues until the performance gain plateaus or a maximum number of epochs is reached. At the end of each cycle, the utility estimates $Q_{i}(l)$ and selection counts $n_{i}(l)$ are updated, informing the sampling policy for the subsequent round. More details are presented in the Supplementary S1.3.
@@
In binary classification, the Area Under the Receiver Operating Characteristic Curve (AUROC) serves as a widely accepted metric for evaluating model performance. We also plot the receiver operating characteristic (ROC) curve to illustrate the diagnostic ability of the classifiers by mapping its true positive rate (model’s ability to correctly identify RBPs) against the false positive rate (frequency of errors in which non-RBPs are misclassified as RBPs). Detail explanation about the calculation can be found in Supplementary S1.4.
@@ 3.1.1 Effect of sampling strategies
First, we investigated the impact of different sampling strategies on model performance. Fig. 3 A indicates that our bandit-based sampling strategy achieves the highest AUROC, particularly in regions with low false positive rates. The w/o Exploration variant represents a degenerate form of the bandit-based strategy where the exploration term in Eq. 1 is removed. While this exploitation-only strategy improves upon uniform sampling, it is consistently outperformed by the full UCB-based approach, highlighting the importance of explicitly encouraging the exploration of under-sampled negatives. Furthermore, we observe that the disparity between sampling strategies is sensitive to the input modality (Fig. S1 in supplementary). Notably, when the model relies only on 1-D sequence features, bandit-based sampling yields an AUROC improvement of over 5%, suggesting that robust negative sampling is particularly critical when structural information is absent.
@@ 3.1.2 Contribution of sequence and structure features
Next, we evaluated the relative contributions of sequence and structure features. As illustrated in Fig. 3 B, the model utilizing only sequence features achieves an AUROC of 0.9213, whereas the structure-only model yields a slightly lower AUROC of 0.9130. This suggests that while structural features are commonly regarded as informative, sequence data may offer greater discriminative power in our identification task. Notably, the integration of both modalities results in the highest performance (AUROC 0.9418), consistently outperforming single-modality approaches. One plausible reason is that the model can utilize the complementary nature of the two features: sequence data provides broad discriminative signals, while structural features help resolve ambiguous cases that are difficult to distinguish via sequence alone. Thus, multimodal integration is essential for maximizing the robustness and accuracy of RBP classification.
@@ 3.1.3 Effectiveness of feature fusion strategies
Finally, we evaluated the effectiveness of different strategies for fusing sequence and structure representations. As shown in Fig. 3 C, simple feature concatenation (Concat) yields the lowest performance (AUROC 0.9146), indicating that simply stacking features offers limited gains. While additive fusion (Sum) provides a slight improvement (AUROC 0.9211), these naive aggregation methods remain insufficient for fully capturing cross-modal complementarity. In contrast, more advanced strategies consistently enhance performance: gated additive fusion increases the AUROC to 0.9314 by adaptively reweighting features, while low-rank multiplicative fusion achieves an AUROC of 0.9265, highlighting the benefits of modeling nonlinear cross-modal interactions. Ultimately, our proposed AEFM method achieves the highest AUROC of 0.9418. By dynamically combining additive and multiplicative interactions, AEFM integrates sequence and structural information more effectively than any single fusion strategy.
@@ 3.2 Comparison with benchmark methods
To further evaluate the predictive performance of SeekRBP, we benchmarked it against several state-of-the-art approaches, encompassing both machine learning-based tools (PhANNs and PhageRBPdetection) and homology-based methods (BLASTp and Pharokka). For the learning-based models, ROC curves were generated using the output prediction probabilities, whereas alignment bit-scores were utilized for the homology-based methods. As illustrated in Fig. 4, SeekRBP achieves the highest AUROC of 0.9418, consistently outperforming all baseline methods. Notably, the homology-based methods, Pharokka and BLASTp, demonstrate strong performance with AUROC scores of 0.9024 and 0.8976, respectively, surpassing the existing learning-based tools. However, the ROC curve for SeekRBP strictly dominates the other methods across the entire threshold spectrum. This performance suggests that our bandit-based sampling strategy effectively identifies highly informative negative examples, enabling the model to learn more discriminative features of RBPs without sacrificing generalization to divergent RBPs.
Figure 4: ROC curves evaluating the predictive performance of SeekRBP (denoted as "Ours") alongside existing machine learning-based tools (PhANNs, PhageRBPdetection) and homology-based methods (BLASTp, Pharokka)
@@ 3.3 Case study: application on vibrio phages
To evaluate the biological relevance and generalizability of our proposed framework, we applied SeekRBP to an independent experimental dataset of Vibrio phages. This dataset, generated by our collaborator Zhang et al. (Zhang et al., 2025), involved phage purification, culturing, and lytic induction, with RBPs subsequently identified through the manual inspection of 3D structural domains. While an absolute ground truth for all RBPs is difficult to establish, we utilized this manually curated set as our reference standard. In this analysis, we first used SeekRBP to identify putative RBPs and then compared their downstream utility for host prediction against the manually curated set.
@@
As shown in Fig. 5, SeekRBP identifies a broader set of RBPs than the manually curated reference, while still recovering a substantial fraction of the known RBPs across various phage genomes. Beyond count-based comparisons, we conducted a structure-based analysis using TM-scores to evaluate the structural plausibility of the newly predicted RBPs. Specifically, we compared the TM-scores of manually selected RBPs missed by our framework (the Observed set) against those identified by SeekRBP but absent from manual annotations (the Predicted set). A TM-score above 0.5 generally indicates that two proteins share the same global fold. As illustrated in Supplementary Fig. S2, while the predicted RBPs exhibit a wider variance in TM-scores, the majority exceed the 0.5 threshold. This indicates that SeekRBP successfully detects divergent RBP candidates with valid structural folds, potentially capturing functional proteins that were overlooked during manual curation.
Figure 5: Comparison of the number of predicted RBPs and manually observed RBPs across individual Vibrio phages. The x-axis denotes different Vibrio phages, and the y-axis represents the number of RBPs. The model predicts more RBPs than those detected manually, reflecting the potential complexity of receptor-binding mechanisms and the limitations of condition-specific assays.
Figure 6: Docking quality (interface pLDDT) between RBPs and outer membrane proteins from host genomes estimated by AlphaFold3. The Predicted set exhibits higher interface pLDDT scores than the Observed set, indicating more stable binding interfaces.
@@
Furthermore, we investigated whether these newly predicted RBPs could enhance downstream host recognition. We compared the utility of the Observed and Predicted sets using two complementary validation approaches. First, we evaluated their performance in a learning-based phage-host interaction prediction task. Using ESM-2 embeddings and a standard multilayer perceptron classifier within a 10-fold stratified cross-validation framework (Supplementary S1.5), the Predicted set yielded a higher mean AUROC (0.737) compared to the Observed set (0.713). Additionally, the standard deviation decreased from 0.099 to 0.073, indicating improved predictive stability (Supplementary Fig. S3). Second, to determine if these predictive gains reflect biological insights, we employed AlphaFold3 to estimate the docking quality, measured by interface pLDDT, between the RBPs and host outer membrane proteins (Supplementary S1.6). Consistent with the classification results, Fig. 6 reveals that the Predicted set exhibits higher interface pLDDT scores, suggesting the formation of more stable binding interfaces. Together, these findings demonstrate that the additional RBPs identified by SeekRBP provide robust biological signals rather than computational noise, effectively supplementing sparse experimental annotations.